12. If a sample of water has a density of 1.025 g/cm, how many liters is the sample with a mass of 2.04 kg?
Answer #1

(12) It is given that, Density (d) = 1.025 g/cm3 and mass (m) = 2.04 kg Since, 1 mL = 1 cm3 So, 1 cm3 volume gives 1.025 g oror 1.025 kg gives 1 L Then 2.04 kg = (1/1.025) x 2.04 L = 1.99 L So, the volume will be 1.99 L.